y=8sin(4x-3pie)-4
graph

Answer by lwsshak3(11628)

y=8sin(4x-3pie)-4
graph
**
I don't have the means to draw a graph, but I can provide you with the information you need to draw one.
..
Standard form of equation for sin function: y=Asin(Bx-C), A=amplitude, Period=2π/B,
Phase shift=C/B.
For given equation: y=8sin(4x-3π)-4
amplitude=8
B=4
Period=2π/B=2π/4=π/2
1/4 period=π/8
Phase Shift=C/B=3π/4 (To the right)
..
Graphing for one period:
Basically, you will be plotting a sin curve with a period of π/2 and amplitude=8, shifting it 3π/4 to the right, then dropping the entire curve 4 units down.